MICROSOLVATION OF BUILDING BLOCKS

V. VAQUERO, I. PENA, S. MATA, \underline{J. C. LÓPEZ}, and J. L. ALONSO, Grupo de Espectroscopía Molecular (GEM), Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47005 Valladolid, Spain.

Laser ablation molecular beam Fourier transform microwave spectroscopy LA-MB-FTMW has provided the first observation of the rotational spectra of the glycine-H_{2}O complex.^{a} As a continuation of this work we present the study of monohydrated complexes of alanine, uracil and thymine and the first observations of the dihydrated clusters glycine- (H_{2}O)_{2} and alanine-(H_{2}O)_{2}. For the microsolvated amino acids only the conformer with cis-COOH configuration and bifurcated NH_{2}⋯O=C H-bond has been observed. Both alanine-H_{2}O and glycine-H_{2}O have similar shapes: the water molecule being bonded through two O-H⋯O hydrogen bonds to the carboxylic group of the amino acid. The dihydrated clusters have also comparable structures where the two water molecules form a cycle with the COOH group through sequential H-bonds. The rotational spectra of uracyl-H_{2}O and thymine-H_{2}O show the same complexity than those of the bare molecules^{b}^{,}^{c} due to the hyperfine structure of two quadrupolar ^{14}N atoms. The detected conformers show comparable structures with the water molecule bonded through two N-H⋯O and O-H⋯O=C hydrogen bonds to the N bases. The structure of hydrogen bond has been investigated by isotopic substitution.

aJ.L. Alonso, E.J. Cocinero, A. Lesarri, M.E. Sanz and J.C. López, Angew. Chem. 45, 3471 (2006)

bV. Vaquero, M.E. Sanz, J.C. López and J.L. Alonso,J. Phys. Chem. A111, 3443 (2007)

cJ.C. López, M.I. Peña, M.E. Sanz and J.L. Alonso, J. Chem. Phys. 126, 191103 (2007)